OpenStudy (gokuporter):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (anonymous):

Hey Gokuporter! One way you could solve this is plug and chug. If you can solve for the left you will see the answer is 75. Now you have to figure out which X value makes this equation true.

TheSmartOne (thesmartone):

a(b+c)=ab+ac so 5(8+7)=??

OpenStudy (gokuporter):

Thanks

TheSmartOne (thesmartone):

Ok No prob.. I guess lol

OpenStudy (anonymous):

75 = 5 x + 35

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

Correct.